### 13 Mar 2026 v1

# Vertex-minor universality of a random graph

Abstract.
Given a graph $G$ and a vertex $v\in V(G)$, a local complementation at $v$ on $G$ is an operation that replaces the induced graph on the neighborhood of $v$ by its complement. A graph $H$ is a vertex-minor if $H$ can be obtained from $G$ by a sequence of vertex deletions and local complementation. A graph is said to be $k$-vertex-minor universal if it contains every $k$-vertex graph on any $k$-subset of vertices as a vertex minor. Previously, Ascoli–Fredrickson–Fredrickson–McFarland–Post [undef] proved that with high probability $G(n,1/2)$ is $\Omega(\sqrt{n})$-vertex-minor universal. Furthermore, they conjectured that with high probability $G(n,p)$ and G(n,1p)G(n,1-p) are $\Omega(p\sqrt{n})$-vertex-minor universal for all $\omega(1/\sqrt{n})\leqslant p\leqslant 1/2$. In this short note, we confirm this conjecture up to an extra logarithm factor and show that this is true with probability $1-2^{-\Omega(p^{2}n)}$ if $\Omega(\log n/\sqrt{n})\leqslant p\leqslant 1/2$.
Department of Mathematics, Massachusetts Institute of Technology. Email: {twchao,zixuanxu}@mit.edu
1. Introduction
Let $G=(V,E)$ be a graph. For a vertex $v\in V$, a local complementation at $v$ is an operation where we replace the neighborhood of $v$ by its complement. Let GvG*v denote the resulting graph after performing local complementation at $v$. We say that two graphs are locally equivalent if one can be obtained from the other by a sequence of local complementations. For a graph $H$ with vertex set $V(H)\subseteq V$, we say that $H$ is a vertex-minor of $G$ if $H$ can be obtained from $G$ via a sequence of vertex deletions and local complementations. Equivalently, $H$ is a vertex-minor of $G$ if $H$ is the induced subgraph on $V(H)$ of a graph locally equivalent to $G$. Since their introduction in works of Bouchet [undefa] and Fon-Der-Flaass [undefd] in the 1980s, vertex-minors have been extensively studied in structural graph theory. In particular, local complementations preserve the cut-rank function of a graph. Thus, vertex-minors have close connections to the study of circle graphs, forbidden minors and related decomposition questions. See the survey on vertex-minors by Kim and Oum [undefe] for a more detailed overview of vertex-minors and related concepts.
Motivated by the connections between vertex-minors and quantum communication, a recent line of work considers the universality question for vertex-minors. A graph $G$ is $k$-vertex-minor universal if for every vertex subset $U\subseteq V(G)$ of size $k$, every graph on $U$ is a vertex-minor of $G$. Note that this is a strong condition where every graph on $k$ vertices must be realized on any specified subset of $k$ vertices in $G$. In this direction, Claudet, Mhalla, and Perdrix [undefc] first showed that there exists $k$-vertex-minor universal graphs with $O(k^{4}\log k)$ vertices and any $k$-vertex-minor universal graph must contain at least $(k-2)^{2}/(2\log_{2}3)$ vertices. Then, Cautrés, Claudet, Mhalla, Perdrix, Savin, and Thomassé [undefb] further showed the existence of $k$-vertex-minor universal graphs with $(2+o(1))k^{2}$ vertices. Most recently, Ascoli, Fredrickson, Fredrickson, McFarland, and Post [undef] showed that almost all graphs on $n$ vertices are $O(\sqrt{n})$-vertex-minor universal by proving the following statement. Recall that $G(n,p)$ denotes the Erdős-Renyí random graph on $n$ vertices with each edge present with probability $p$.
Theorem 1.1 ([undef]).
For a positive integer $k$ and $G\sim G(n,1/2)$, if $n\geqslant(1+C)\frac{1}{2\log_{2}(4/3)}k^{2}$ for some $C>0$, then $G$ is $k$-vertex-minor universal with probability at least $1-2^{-(1+o(1))Ck^{2}/2}$.
Furthermore, the authors conjecture that an analogous result holds for $G(n,p)$ and G(n,1p)G(n,1-p) where $p=\omega(1/\sqrt{n})$ and $p\leqslant 1/2$. Formally, they conjecture the following statement.
Conjecture 1.2 ([undef]).
If $p=\omega(1/\sqrt{n})$ and $p\leqslant 1/2$ and $G\sim G(n,p)$ or $G\sim G(n,1-p)$, then $G$ is $k$-vertex-minor universal with high probability for some $k=\Omega(p\sqrt{n})$.
It was remarked in [undef] that the method used to prove Theorem 1.1 cannot be naturally extended to other values of $p\neq 1/2$. This is because Theorem 1.1 is proved by considering the random walk on the set of graphs on a fixed subset $U$ of size $k$ produced by a random sequence of local complementations at vertices outside of $U$. A crucial property used in the proof is the fact that each step of the random walk is independent when $p=1/2$, which no longer holds for $p\neq 1/2$. Our main result is to confirm 1.2 up to an extra logarithm factor.
Theorem 1.3.
There exists a constant $n_{0}\in\mathbb{N}$ such that the following holds for all $n>n_{0}$. Suppose that $p\in[0,1]$ and $q\overset{\rm{def}}{=}\min(p,1-p)\geqslant 100\log n/\sqrt{n}$. For any $k\leqslant q\sqrt{n}/100$, and $G\sim G(n,p)$, the probability that $G$ is $k$-vertex-minor universal is at least $1-2^{-q^{2}n/100}$.
Our main technical contribution is Lemma 3.1, which formalizes the following intuition: Let $G\sim G(n,p)$ and fix a set of $\Omega(k^{2})$ vertices $U\subseteq V(G)$. Let $W\overset{\rm{def}}{=}V(G)\setminus U$. Then after a sequence of local complementations at $\{w_{1},\dots,w_{r}\}=W$, the distribution of the induced subgraph $(G*w_{1}*\dots*w_{r})[U]$ should be close to $G(U,1/2)$ in total variation distance. Indeed, if $W$ is an independent set, any edge $uv$ with $u,v\in U$ is toggled during local complementation at $w_{i}$ if $u,v\in N(w_{i})$. Furthermore, since $W$ is an independent set, local complementation at $w_{j}$ for $j<i$ does not toggle the edges between $w_{i}$ and $U$. So, the event that $u,v\in N(w_{i})$ holds with probability $p^{2}$ and only depends on the edges between $w_{i}$ and $U$. Moreover, the events that $u,v\in N(w_{i})$ are independent for all $w_{i}$ where $i\in[r]$. Therefore, by a direct calculation, the probability that $uv$ is an edge in $(G*w_{1}*\dots*w_{r})[U]$ is $\frac{1}{2}\pm 2^{-\Omega(q^{2}r)}$. The main difficulty is to generalize this observation from the distribution over a single edge $uv$ to the distribution over all possible graphs on $U$, and to deal with the dependencies between local complementations at $w_{1},\dots,w_{r}$ when $W$ is not an independent set in $G$. Once we have proved Lemma 3.1, we can prove Theorem 1.3 by applying Theorem 1.1 to $(G*w_{1}*\dots*w_{r})[U]$ in a black-box way.
Notations. For a graph $G$ and a subset $U\subseteq V(G)$, we use $G[U]$ to denote the induced subgraph of $G$ on $U$. For vertex sets $A,B\subseteq V(G)$, we use $G[A,B]$ to denote the induced bipartite graph of $G$ on $A\times B$. For sets $A$ and $B$, we denote their symmetric difference as $A\triangle B\overset{\rm{def}}{=}(A\setminus B)\cup(B\setminus A)$. For $p\in[0,1]$, we use $x\sim\textup{Ber}(p)$ to denote that $x$ is a Bernoulli random variable taking value $1$ with probability $p$ and 0 with probability 1p1-p. For any matrix $A$, we use $A_{ij}$ to denote the $(i,j)$-entry of $A$. For an $m\times n$ matrix $A$ and a $m^{\prime}\times n^{\prime}$ matrix $B$, the tensor product $A\otimes B$ is the $(mm^{\prime})\times(nn^{\prime})$ matrix indexed by pairs $(i,i^{\prime})\in[m]\times[m^{\prime}]$ and $(j,j^{\prime})\in[n]\times[n^{\prime}]$ with entries $(A\otimes B)_{(i,i^{\prime}),(j,j^{\prime})}=A_{ij}B_{i^{\prime}j^{\prime}}$. For a matrix $M\in\mathbb{F}_{2}^{m\times n}$, we use $\textup{rank}(M)$ to denote its rank over $\mathbb{F}_{2}$.
Paper Organization. In Section 2, we present a preliminary lemma that will be used in proving Theorem 1.3. Then in Section 3, we give the proof of Theorem 1.3. Finally, in Section 4, we include some concluding remarks about applying our techniques to other related settings.
2. Preliminary
In this section, we present useful definitions and lemmas that will be used in our proof of Theorem 1.3. We start with presenting a useful lemma that gives an upper bound for $\left|\mathbb{E}[(-1)^{f(x_{1},\dots,x_{m})}\right|$ where $x_{1},\dots,x_{m}$ are independent $\textup{Ber}(p)$ random variables and $f(x_{1},\dots,x_{m})$ is a multilinear polynomial of degree at most $2$ over $\mathbb{F}_{2}$ in terms of the rank of the coefficient matrix of degree $2$ terms in $f$. For any multilinear polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ with $\deg f\leqslant 2$, we use $\textup{Coef}_{2}(f)$ to denote the $m\times m$ coefficient matrix over $\mathbb{F}_{2}$ with entries $(a_{ij})_{i,j=1}^{m}$ that records the degree $2$ coefficient of $f$. Namely, $a_{ij}=1$ if $x_{i}x_{j}$ is a monomial in $f$ and $a_{ij}=0$ otherwise. In particular, since $f$ is multilinear, every entry on the diagonal of $\textup{Coef}_{2}(f)$ is 0.
Lemma 2.1.
Let $p\in[0,1]$ and $q\overset{\rm{def}}{=}\min(p,1-p)$. Let $r$ be a positive integer and $X\in\mathbb{F}_{2}^{m}$ be a vector with independent $\textup{Ber}(p)$ entries. Let $f\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ be a multilinear polynomial with $\deg f\leqslant 2$. Suppose that $\textup{rank}(\textup{Coef}_{2}(f))=r$. Then we have
$\left|\mathbb{E}[(-1)^{f(X)}]\right|\leqslant(1-2q^{2})^{\lfloor r/6\rfloor}.$
Proof.
We prove by induction on $r$. In the base case where $r\leqslant 6$, the statement is trivial since the right hand side is $1$.
Now suppose the statement is true for all integers less than $r$ and $r\geqslant 6$. Let $X=(X_{1},\dots,X_{m})\in\mathbb{F}_{2}^{m}$ be a vector with independent $\textup{Ber}(p)$ entries. Let $f\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ be a multilinear polynomial with $\deg f\leqslant 2$ such that $\textup{rank}(\textup{Coef}_{2}(f))=r$. Let $\textup{Coef}_{2}(f)=(a_{ij})_{i,j=1}^{m}$ and we can further assume without loss of generality that $a_{12}=1$. Then we can write $f$ as
$f=x_{1}x_{2}+x_{1}f_{1}+x_{2}f_{2}+f_{3},$
where $f_{1},f_{2},f_{3}\in\mathbb{F}_{2}[x_{3},\dots,x_{m}]$ such that $\deg f_{1},\deg f_{2}\leqslant 1$, $\deg f_{3}\leqslant 2$, and $f_{3}$ is multilinear. Note that $\textup{Coef}_{2}(f_{3})$ is a minor of $\textup{Coef}_{2}(f)$ obtained from removing the first two rows and columns from $\textup{Coef}_{2}(f)$. Therefore, we have $\textup{rank}(\textup{Coef}_{2}(f_{3}))\geqslant r-4$. Denote $X^{\prime}=(X_{3},\dots,X_{m})$ and we have
$\mathbb{E}[(-1)^{f(X)}]=\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})}\mathbb{E}_{X_{1},X_{2}}[(-1)^{X_{1}X_{2}+X_{1}f_{1}(X^{\prime})+X_{2}f_{2}(X^{\prime})}\mid X^{\prime}]\right].$
We first prove the case $q=p\leqslant 1/2$. By a direct calculation, for any $a,b\in\mathbb{F}_{2}$ and since $X_{1},X_{2}\sim\textup{Ber}(p)$ are independent, we have
$\mathbb{E}_{X_{1},X_{2}}[(-1)^{X_{1}X_{2}+aX_{1}+bX_{2}}]=(1-p)^{2}+p(1-p)((-1)^{a}+(-1)^{b})-p^{2}(-1)^{a+b}.$ (1)
Using the identity that
$(-1)^{a+b}=1+((-1)^{a}+(-1)^{b})-2(-1)^{ab},$
we obtain
$\mathbb{E}_{X_{1},X_{2}}[(-1)^{X_{1}X_{2}+aX_{1}+bX_{2}}]=(1-2p)+p(1-2p)((-1)^{a}+(-1)^{b})+2p^{2}(-1)^{ab}.$ (2)
Therefore, we can write
$\mathbb{E}[(-1)^{f(X)}]=$ $\,\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})}\left((1-2p)+p(1-2p)((-1)^{f_{1}(X^{\prime})}+(-1)^{f_{2}(X^{\prime})})+2p^{2}(-1)^{f_{1}(X^{\prime})f_{2}(X^{\prime})}\right)\right]$
$=$ $\,(1-2p)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})}\right]+p(1-2p)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})}\right]$
$+p(1-2p)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{2}(X^{\prime})}\right]+2p^{2}\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})f_{2}(X^{\prime})}\right].$
Now, we can apply the inductive hypothesis to bound each term. Recall that we have $\textup{rank}(\textup{Coef}_{2}(f_{3}))\geqslant r-4$. Therefore, we obtain
$\left|\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})}\right]\right|\leqslant(1-2p^{2})^{\lfloor(r-4)/6\rfloor}\leqslant(1-2p^{2})^{\lfloor r/6\rfloor-1}.$
Furthermore, since $\deg f_{1},\deg f_{2}\leqslant 1$, we have $\textup{Coef}_{2}(f_{3}+f_{1})=\textup{Coef}_{2}(f_{3}+f_{2})=\textup{Coef}_{2}(f_{3})$, which has rank at least r4r-4. Thus, we can conclude similarly that
$\left|\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})}\right]\right|\leqslant(1-2p^{2})^{\lfloor r/6\rfloor-1}\text{ and }\left|\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{2}(X^{\prime})}\right]\right|\leqslant(1-2p^{2})^{\lfloor r/6\rfloor-1}.$
Finally, note that the homogeneous degree $2$ part of $f_{1}f_{2}$ is $\sum_{i,j=3}^{m}\alpha_{i}\beta_{j}x_{i}x_{j}$ if $f_{1}=\sum_{i=3}^{m}\alpha_{i}x_{i}+c_{1}$, and $f_{2}=\sum_{j=3}^{m}\beta_{j}x_{j}+c_{2}$. Consider the multilinearization $g$ of $f_{1}f_{2}$, i.e. the polynomial $g$ obtained by replacing all the terms $x_{i}^{2}$ by $x_{i}$ in $f_{1}f_{2}$. It follows that the coefficient of $x_{i}x_{j}$ in $g$ is $\alpha_{i}\beta_{j}+\alpha_{j}\beta_{i}$ if $i\neq j$. Therefore, we have
$\textup{Coef}_{2}(g)=\alpha^{T}\beta+\beta^{T}\alpha$
where $\alpha=(\alpha_{3},\dots,\alpha_{m})$ and $\beta=(\beta_{3},\dots,\beta_{m})$. Since the diagonal terms of $\alpha^{T}\beta+\beta^{T}\alpha$ are always 0, they agree with the diagonal terms of $\textup{Coef}_{2}(g)$. Therefore, we have $\textup{rank}(\textup{Coef}_{2}(g))\leqslant 2$. Hence, we obtain that $\textup{rank}(\textup{Coef}_{2}(f_{3}+g))\geqslant r-6$. By the inductive hypothesis, we have
$\left|\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})f_{2}(X^{\prime})}\right]\right|=\left|\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+g(X^{\prime})}\right]\right|\leqslant(1-2p^{2})^{\lfloor(r-6)/6\rfloor}=(1-2p^{2})^{\lfloor r/6\rfloor-1}.$
Combining the above inequalities and the fact that $p,(1-2p)\geqslant 0$ for $p\leqslant 1/2$, we obtain
$\left|\mathbb{E}[(-1)^{f(X)}]\right|\leqslant\left((1-2p)+p(1-2p)+p(1-2p)+2p^{2}\right)(1-2p^{2})^{\lfloor r/6\rfloor-1}=(1-2p^{2})^{\lfloor r/6\rfloor}.$
For the case $p\geqslant 1/2$ and q=1pq=1-p, we can combine Eq. 1 with the identity
$1=(-1)^{a+b}-((-1)^{a}+(-1)^{b})+2(-1)^{ab}$
and conclude that
$\mathbb{E}_{X_{1},X_{2}}[(-1)^{X_{1}X_{2}+aX_{1}+bX_{2}}]=$ $-(2p-1)(-1)^{a+b}+(1-p)(2p-1)((-1)^{a}+(-1)^{b})+2(1-p)^{2}(-1)^{ab}$
$=$ $-(1-2q)(-1)^{a+b}+q(1-2q)((-1)^{a}+(-1)^{b})+2q^{2}(-1)^{ab}.$
Similar as before, we have
$\mathbb{E}[(-1)^{f(X)}]$
$=\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})}\left(-(1-2q)(-1)^{f_{1}(X^{\prime})+f_{2}(X^{\prime})}+q(1-2q)((-1)^{f_{1}(X^{\prime})}+(-1)^{f_{2}(X^{\prime})})+2q^{2}(-1)^{f_{1}(X^{\prime})f_{2}(X^{\prime})}\right)\right]$
$=-(1-2q)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})+f_{2}(X^{\prime})}\right]+q(1-2q)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})}\right]$
$\quad+q(1-2q)\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{2}(X^{\prime})}\right]+2q^{2}\mathbb{E}_{X^{\prime}}\left[(-1)^{f_{3}(X^{\prime})+f_{1}(X^{\prime})f_{2}(X^{\prime})}\right].$
$\leqslant\left((1-2q)+q(1-2q)+q(1-2q)+2q^{2}\right)(1-2q^{2})^{\lfloor r/6\rfloor-1}=(1-2q^{2})^{\lfloor r/6\rfloor},$
where the last inequality follows from applying the inductive hypothesis to $f_{3}+f_{1}+f_{2}$, $f_{3}+f_{1}$, $f_{3}+f_{2}$, and $f_{3}+g$, where $g$ is the multilinearization of $f_{1}f_{2}$. We also used the fact that $q,(1-2q)\geqslant 0$. This concludes the proof. ∎
In the proof of Lemma 3.1, we will also use the following upper bound on the number of graphs that has adjacency matrix of a fixed rank over $\mathbb{F}_{2}$. For a graph $F$, we use $\textup{rank}(F)$ to denote the rank of its adjacency matrix over $\mathbb{F}_{2}$.
Theorem 2.2.
[undeff, Theorem 3] For a positive integer $s$ and an integer $a\geqslant 0$, the number of graphs $F$ on the vertex set $[s]$ such that $\textup{rank}(F)=a$ is at most $2^{sa-2}$.
Remark 2.3.
In [undeff, Theorem 3], it was shown that if $a$ is odd, then the number of graphs $F$ with $\textup{rank}(F)=a$ is 0. If $a=2b$ is even, then the number of graphs $F$ with $\textup{rank}(F)=a$ is exactly
$\prod_{i=1}^{b}\frac{2^{2i-2}}{2^{2i}-1}\prod_{i=0}^{2b-1}(2^{s-i}-1).$
Note that we have $\prod_{i=1}^{b}2^{2i-2}=2^{b^{2}-b}$, $\prod_{i=1}^{b}(2^{2i}-1)\geqslant\prod_{i=1}^{b}2^{2i-1}=2^{b^{2}}$, and
$\prod_{i=0}^{2b-1}(2^{s-i}-1)\leqslant\prod_{i=0}^{2b-1}2^{s-i}=2^{2sb-2b^{2}+b}.$
Thus, we have
$\prod_{i=1}^{b}\frac{2^{2i-2}}{2^{2i}-1}\prod_{i=0}^{2b-1}(2^{s-i}-1)\leqslant 2^{2sb-2b^{2}}\leqslant 2^{sa-2},$
which is the upper bound stated in Theorem 2.2.
3. Proof of Theorem 1.3
[…]
Lemma 3.1.
For $p\in[0,1]$, $q\overset{\rm{def}}{=}\min(p,1-p)$, and $G\sim G(n,p)$. For any vertex partition $V(G)=U\sqcup W$ with $\left|U\right|=s$ and $\left|W\right|=r$ satisfying $s\leqslant q^{2}r/6$. For any ordering on the vertices of $W=\{w_{1},\dots,w_{r}\}$, let $G^{\prime}=G*w_{1}*\dots*w_{r}$. Then the total variation distance between $G^{\prime}[U]$ and $G(U,1/2)$ is at most $2^{-q^{2}r/6}$.
Proof.
Let $G_{\Delta}\overset{\rm{def}}{=}G^{\prime}[U]\triangle G[U]$. Note that since $G_{\Delta}$ only depends on $G[U,W]$ and $G[W]$, it is independent from $G[U]$. Moreover, note that the symmetric difference between $G[U]\sim G(U,p)$ and an independent copy of $G(U,1/2)$ has the same distribution as $G(U,1/2)$. Indeed, this is because the indicator of each edge $e$ in the symmetric difference is of the form $X_{e}+Y_{e}$ where $X_{e}\sim\textup{Ber}(p)$ and $Y_{e}\sim\textup{Ber}(1/2)$ are independent, so $X_{e}+Y_{e}\sim\textup{Ber}(1/2)$. Thus, it suffices to show that the total variation distance between $G_{\Delta}$ and $G(U,1/2)$ is at most $2^{-q^{2}r/6}$.
Label the vertices in $U$ as $\{u_{1},\dots,u_{s}\}=U$ and let $X\in\mathbb{F}_{2}^{s\times r}$ be the $s\times r$ matrix with entries $X_{ij}$ being the indicator variable of $u_{i}w_{j}\in E(G)$ for all $i\in[s]$ and $j\in[r]$. Note that by definition, $X$ has independent $\textup{Ber}(p)$ entries corresponding to the biadjacency matrix of $G[U,W]$. For a square matrix $M$, we use $\textup{OffDiag}(M)\overset{\rm{def}}{=}M-\textup{Diag}(M)$ to denote the matrix obtained by setting all the diagonal entries of $M$ to zero.
Claim 3.2.
There is an invertible symmetric $r\times r$ matrix $M\in\mathbb{F}_{2}^{r\times r}$ that only depends on $G[W]$ such that $\textup{OffDiag}(XMX^{T})$ is the adjacency matrix of $G_{\Delta}$.
Proof.
For each $i\in[r]$, let $G_{i}\overset{\rm{def}}{=}G*w_{1}*\dots*w_{i}$. Let $z_{i}\in\mathbb{F}_{2}^{s}$ be the indicator vector of the edges in $G_{i}[U,w_{i}]$ and let $X_{i}$ be the $i$-th column vector of $X$. Note that $z_{i}=X_{i}+\sum_{j}z_{j}$, where the summation is over all $j<i$ such that $w_{i}w_{j}$ is an edge in $G_{j-1}$. This is because for each $j<i$, local complementation at $w_{j}$ on $G_{j-1}$ flips the edge between $w_{i}u$ for $u\in U$ if and only if $w_{i}w_{j}$ and $w_{j}u$ are both edges in $G_{j-1}$. Consider the matrix $L\in\mathbb{F}_{2}^{r\times r}$ defined as
[…]
Furthermore, note that $L$ is upper triangular with all diagonal entries being $1$, so $L$ is invertible. Then $XL^{-1}$ is the matrix with columns exactly $(z_{i})_{i=1}^{r}$. Now, we claim that $M=L^{-1}(L^{-1})^{T}$ satisfies the conditions in the claim. First, it is clear that $L$ only depends on $G_{j-1}[W]$, which only depends on $G[W]$. Thus, $M$ only depends on $G[W]$.
Note that from the definition of local complementation, the edge $u_{j}u_{k}$ is flipped during local complementation at $w_{i}$ on $G_{i-1}$ if and only if both $u_{j}w_{i}$ and $u_{k}w_{i}$ are edges in $G_{i-1}$. Thus, the adjacency matrices of $G_{i-1}[U]$ and $G_{i}[U]$ differ by exactly $\textup{OffDiag}(z_{i}z_{i}^{T})$. It follows directly that the adjacency matrix of $G_{\Delta}$ is exactly $\textup{OffDiag}(\sum_{i=1}^{r}z_{i}z_{i}^{T})=\textup{OffDiag}(XMX^{T})$. This completes the proof. ∎
Next, we analyze the distribution of $G_{\Delta}$ using discrete Fourier transformation. To simplify the analysis, we will always condition on $G[W]=G^{*}$ from now on. We use $\mu:\mathbb{F}_{2}^{\binom{U}{2}}\rightarrow[0,1]$ to denote the probability distribution of $G_{\Delta}$ conditioning on $G[W]=G^{*}$. Namely, for each $H\subseteq\binom{U}{2}$, we define
$\mu(H)\overset{\rm{def}}{=}\mathbb{P}(G_{\Delta}=H\mid G[W]=G^{*}).$
In addition, let $\mu_{\textup{unif}}$ be the uniform distribution over $\binom{U}{2}$, i.e. $\mu_{\textup{unif}}=2^{-\binom{s}{2}}$ where $s=|U|$ for every $H\subseteq\binom{U}{2}$.
For a function $f:\mathbb{F}_{2}^{\binom{U}{2}}\rightarrow\mathbb{R}$, we define its Fourier transformation $\widehat{f}:\mathbb{F}_{2}^{\binom{U}{2}}\rightarrow\mathbb{R}$ by
$\widehat{f}(F)\overset{\rm{def}}{=}\sum_{H\subseteq\binom{U}{2}}f(H)(-1)^{\left|H\cap F\right|}.$
Here we abuse notation and use $H$ and $F$ to denote the graphs on the vertex set $U$ and their corresponding edge indicator vector in $\mathbb{F}_{2}^{\binom{U}{2}}$. We first recall the following standard fact of bounding the total variation distance between $\mu$ and $\mu_{\textup{unif}}$ via the Fourier coefficients of $\mu$.
Claim 3.3.
Let $\mu$ be a probability distribution over $\binom{U}{2}$ and $\mu_{\textup{unif}}$ be the uniform distribution over $\binom{U}{2}$. Then we have
$d_{\textup{TV}}(\mu,\mu_{\textup{unif}})\leqslant\frac{1}{2}\sum_{F\neq\varnothing\subseteq\binom{U}{2}}|\widehat{\mu}(F)|.$
Proof.
Note that we have
$\widehat{\mu_{\textup{unif}}}(F)=\begin{cases}1&\text{ if }F=\varnothing,\\ 0&\text{ otherwise},\end{cases}$
and $\widehat{\mu}(\varnothing)=1$. Thus, by linearity of the Fourier transform, we have
$\widehat{(\mu-\mu_{\textup{unif}})}(F)=\widehat{\mu}(F)-\widehat{\mu_{\textup{unif}}}(F)=\begin{cases}0&\text{ if }F=\varnothing,\\ \widehat{\mu}(F)&\text{ otherwise}.\end{cases}$
[…]
$\left|f(H)\right|\leqslant 2^{-\binom{s}{2}}\sum_{F\subseteq\binom{U}{2}}\left|\widehat{f}(F)\right|.$
Therefore,
$f||_{1}=\sum_{H\subseteq\binom{U}{2}}\left|f(H)\right|\leqslant 2^{-\binom{s}{2}}\sum_{H\subseteq\binom{U}{2}}\sum_{F\subseteq\binom{U}{2}}\left|\widehat{f}(F)\right|=\sum_{F\subseteq\binom{U}{2}}\left|\widehat{f}(F)\right|.\qed$
Now we upper bound $\left|\widehat{\mu}(F)\right|$ for each $F\subseteq\binom{U}{2}$ by the rank of its adjacency matrix. From now on, we abuse notation and use $\textup{rank}(F)$ to denote the rank of the adjacency matrix of $F$ over $\mathbb{F}_{2}$.
Claim 3.4.
For any $F\subseteq\binom{U}{2}$, we have
$\left|\widehat{\mu}(F)\right|\leqslant(1-2q^{2})^{r\cdot\textup{rank}(F)/6-1}.$
Proof.
Recall that by definition, we have
$\widehat{\mu}(F)=\sum_{H\subseteq\binom{U}{2}}\mathbb{P}(G_{\Delta}=H\mid G[W]=G^{*})(-1)^{\left|H\cap F\right|}=\mathbb{E}_{H\sim\mu}[(-1)^{\left|H\cap F\right|}],$
where we recall that $\mu$ is the distribution of $(G_{\Delta}\mid G[W]=G^{*})$. By Claim 3.2, we know that there exists an invertible symmetric matrix $M$ satisfying the following: Let $X\in\mathbb{F}_{2}^{s\times r}$ be the biadjacency matrix of $G[U,W]$ over $\mathbb{F}_{2}$ with independent $\textup{Ber}(p)$ entries. Let $v_{i}\overset{\rm{def}}{=}(X_{i1},\dots,X_{ir})$ be the $i$-row of $X$. Then for any $\{i,j\}\in\binom{[s]}{2}$, we have $u_{i}u_{j}\in H$ if and only if $v_{i}Mv_{j}^{T}=1$.
Define for convenience $F_{0}\overset{\rm{def}}{=}\{\{i,j\}\in\binom{[s]}{2}\mid u_{i}u_{j}\in F\}$ so that $F_{0}$ is the graph $F$ relabeled on the vertex set $[s]$ via the ordering $u_{i}\mapsto i$ on $U$. It follows that $\left|H\cap F\right|\equiv\sum_{\{i,j\}\in F_{0}}v_{i}Mv_{j}^{T}\pmod{2}$. Therefore, we can rewrite
$\widehat{\mu}(F)=\mathbb{E}_{X}\left[(-1)^{\sum_{ij\in F_{0}}v_{i}Mv_{j}^{T}}\right].$
Let $f\in\mathbb{F}_{2}[X_{11},\dots,X_{sr}]$ be the degree-$2$ polynomial given by $f(X)=\sum_{\{i,j\}\in F_{0}}v_{i}Mv_{j}^{T}$. Note that $f$ is multilinear since there are no self-loops in $F_{0}$. Furthermore, for any $i,k\in[s]$ and $j,\ell\in[r]$, the term $X_{ij}X_{k\ell}$ appears in $f$ if and only if $\{i,k\}\in F_{0}$ and $M_{j\ell}=1$. Thus, we know that $\textup{Coef}_{2}(f)=A_{F_{0}}\otimes M$, where $A_{F_{0}}$ is the adjacency matrix of $F_{0}$. Namely, the coefficient of the term $X_{ij}X_{k\ell}$ in $f$ is $(A_{F_{0}})_{ik}M_{j\ell}$. Therefore, we have $\textup{rank}(\textup{Coef}_{2}(f))=\textup{rank}(M)\cdot\textup{rank}(F_{0})=r\cdot\textup{rank}(F)$ since $F_{0}$ is just $F$ on the vertex set $[s]$. By Lemma 2.1, we can conclude that
$\left|\widehat{\mu}(F)\right|=\left|\mathbb{E}_{X}\left[(-1)^{f(X)}\right]\right|\leqslant(1-2q^{2})^{\lfloor r\cdot\textup{rank}(F)/6\rfloor}\leqslant(1-2q^{2})^{r\cdot\textup{rank}(F)/6-1}.\qed$
Now for any $a\geqslant 0$, we have by Theorem 1.3 the number of graphs $F\subseteq\binom{U}{2}$ with $\textup{rank}(F)=a$ is at most $2^{sa-2}$. Combined with Claims 3.3 and 3.4, we obtain
$d_{\textup{TV}}(\mu,\mu_{\textup{unif}})\leqslant$ $\frac{1}{2}\sum_{F\neq\varnothing\subseteq\binom{U}{2}}|\widehat{\mu}(F)|\leqslant\frac{1}{2}\sum_{F\neq\varnothing\subseteq\binom{U}{2}}(1-2q^{2})^{r\cdot\textup{rank}(F)/6-1}\leqslant\frac{1}{2}\sum_{a=1}^{\infty}2^{sa-2}(1-2q^{2})^{ra/6-1}$
$\leqslant$ $\frac{1}{2}\sum_{a=1}^{\infty}2^{sa}2^{-2q^{2}ra/6},$
where the last inequality follows the fact that $1-2q^{2}\geqslant 1/4$ and for all $x\geqslant 0$ we have $1-x\leqslant 2^{-x}$. Note that the common ratio in the geometric series is $2^{s-2q^{2}r/6}\leqslant 2^{-s}\leqslant 1/2$ since $s-2q^{2}r/6\leqslant-s\leqslant-1$. Therefore, we have
$\frac{1}{2}\sum_{a=1}^{\infty}2^{sa-2q^{2}ra/6}\leqslant 2^{s-2q^{2}r/6}\leqslant 2^{-q^{2}r/6}.$
By an averaging argument over all possible assignments $G^{*}$ to $G[W]$, we can conclude that
$d_{\textup{TV}}(G_{\Delta},G(U,1/2))\leqslant 2^{-q^{2}r/6}$
as desired. ∎
Now, we can prove Theorem 1.3.
Proof of Theorem 1.3.
It suffices to prove the statement for $k=\lfloor q\sqrt{n}/100\rfloor$ since $k$-vertex-minor universality is monotone decreasing in $k$. Let $G\sim G(n,p)$. For any fixed $S\subseteq[n]$ with size $k$. We partition the vertices $V(G)=U\sqcup W$ with $S\subseteq U$, $\left|U\right|=\lfloor q^{2}n/12\rfloor$, and $W=\{w_{1},\dots,w_{r}\}$ with $r=n-\lfloor q^{2}n/12\rfloor$. When $n$ is sufficiently large, we have $r\geqslant n/2$. Therefore, we have $q^{2}r/6\geqslant q^{2}n/12\geqslant\lfloor q^{2}n/12\rfloor=s$. Let $G^{\prime}=G*w_{1}*\dots*w_{r}$. By Lemma 3.1, we have the total variation distance between $G^{\prime}[U]$ and $G(U,1/2)$ is at most $2^{-q^{2}n/12}$. Therefore, there is a coupling such that
$\mathbb{P}(G^{\prime}[U]\neq G(U,1/2))\leqslant 2^{-q^{2}n/12}.$
Applying Theorem 1.1 to $U$ with $\left|U\right|=\lfloor q^{2}n/12\rfloor$ and pick $C=2\log_{2}(4/3)\lfloor q^{2}n/12\rfloor/k^{2}-1$. When $n$ is sufficiently large, we know that $C>q^{2}n/15k^{2}$, and hence we have
$\mathbb{P}(G(U,1/2)\text{ is not vertex-minor universal})\leqslant 2^{-(1+o(1))Ck^{2}/2}\leqslant 2^{-q^{2}n/40}.$
Since we have
[…]
we can conclude that for $n$ sufficiently large, we have
$\mathbb{P}(G^{\prime}[U]\text{ is not vertex-minor universal})\leqslant 2^{-q^{2}n/12}+2^{-q^{2}n/40}\leqslant 2^{1-q^{2}n/40}.$
For each $S\subseteq[n]$ with size $k$, we fix a partition $V(G)=U\sqcup W$ as above and we have the probability that there is a graph on vertex set $S$ that is not a vertex-minor of $G$ is at most $2^{1-q^{2}n/40}$. Now, we take a union bound over all possible $S$. Since there are $\binom{n}{k}\leqslant n^{k}$ many subsets $S$ and by assumption we have $k=\lfloor q\sqrt{n}/100\rfloor$, the number of subsets $S$ is at most $n^{\lfloor q\sqrt{n}/100\rfloor}\leqslant 2^{q\sqrt{n}\log n}\leqslant 2^{q^{2}n/100}$. Therefore, by the union bound, $G$ is not $k$-vertex-minor universal with probability at most $2^{q^{2}n/100}\cdot 2^{1-q^{2}n/40}\leqslant 2^{-q^{2}n/100}$ for $n$ sufficiently large. This completes the proof. ∎
4. Concluding remarks
In [undef], the authors also considered the universality question for pivot-minors and matroid-minors and proved analogous results to Theorem 1.1 for $G(n,1/2)$ in these settings. We believe our methods should be able to generalize those results to the setting of $G(n,p)$. We briefly outline how to adapt our strategy to those settings and discuss the extra obstruction in this section.
Since minors in binary matroids can be represented by pivot-minors of their fundamental graphs in the ordered bipartite setting, let us follow the setup in [undef, Section 4] and work with ordered bipartite random graph from now on. See [undef] for more details on the reduction from binary matroids to pivot minors. For a simple graph $G$ and an edge $uv\in E(G)$, a pivot at $uv$, denoted by $G\times uv$, is defined to be the operation GuvuG*u*v*u. One can check that Guvu=GvuvG*u*v*u=G*v*u*v, so the pivot operation does not depend on the ordering of $uv$ and thus is well-defined. Equivalently, the graph $G\times uv$ is obtained by taking the symmetric difference between $G$ with the complete tripartite graph with parts $N(v)\cap N(u)$, $N(v)\setminus(N(u)\setminus\{u\})$, and $N(u)\setminus(N(v)\setminus\{v\})$, and then swapping the labels of $u$ and $v$. An ordered bipartite graph $G=(L,R,E)$ has bipartition $V(G)=L\sqcup R$ and edges contained in $L\times R$. Note that for $uv\in E(G)$, we have
$G\times uv\overset{\rm{def}}{=}(L\triangle\{u,v\},R\triangle\{u,v\},E^{\prime})$
where $E^{\prime}$ is the edges of the unordered bipartite graph $(L,R,E)\times uv$. For any ordered bipartite graph $H=(V_{L},V_{R},E(H))$ with $U_{L}\cup U_{R}\subseteq V(G)$, we say that $H$ is a (labeled) pivot-minor of $G$ if it can be obtained on $V_{L}\cup V_{R}$ via a sequence of vertex deletions and pivots. We say that ordered bipartite graph $G$ is $k$-pivot-minor universal if any ordered bipartite graph $H$ on any $k$ vertices of $V(G)$ is a pivot minor of $G$.
Let $G(\ell,r,p)$ denote the random ordered bipartite graph on vertex bipartition $L\sqcup R$ with $|L|=\ell$, $|R|=r$, and edges between $L\times R$ present independently with probability $p$. We believe Lemma 3.1 can be adapted in this setting with the following strategy.
We fix disjoint subsets $U_{L}\sqcup W_{L}\subseteq L$ and $U_{R}\sqcup W_{R}\subseteq R$ such that $|U_{L}|=|U_{R}|=s$ and $|W_{L}|=|W_{R}|=r$. Suppose that we can find disjoint pairs $w_{1}w^{\prime}_{1},\dots,w_{r^{\prime}}w^{\prime}_{r^{\prime}}\in W_{L}\times W_{R}$ such that $w_{i+1}w^{\prime}_{i+1}$ is an edge in $G_{i-1}=G\times w_{1}w^{\prime}_{1}\times\dots\times w_{i}w^{\prime}_{i}$ for all $i=0,1,\dots,r^{\prime}-1$. Then we set $G^{\prime}=G_{r^{\prime}}$ to be the resulting graph, and let $W^{\prime}_{L}=\{w_{1},\dots,w_{r^{\prime}}\},W^{\prime}_{R}=\{w^{\prime}_{1},\dots,w^{\prime}_{r^{\prime}}\}$.
Define $G_{\Delta}=G^{\prime}[U_{L},U_{R}]\triangle G[U_{L},U_{R}]$. Let $G_{i}=G\times w_{1}w^{\prime}_{1}\times\dots\times w_{i}w^{\prime}_{i}$ and $U_{L}=\{u_{1},\dots,u_{s_{1}}\}$, $U_{R}=\{u^{\prime}_{1},\dots,u^{\prime}_{s_{2}}\}$. Then we can define $X_{L}\in\mathbb{F}_{2}^{|U_{L}|\times|W^{\prime}_{R}|}$ with entries $(X_{L})_{ij}$ being the indicator variable of the edge $u_{i}w^{\prime}_{j}$. Similarly, define $X_{R}\in\mathbb{F}_{2}^{|U_{R}|\times|W^{\prime}_{L}|}$ with entries $(X_{R})_{ij}$ being the indicator variable of the edge $u^{\prime}_{i}w_{j}$. Then similar to Claim 3.2, we claim that there is an invertible (not necessarily symmetric) matrix $M\in\mathbb{F}_{2}^{r^{\prime}\times r^{\prime}}$ only depending on $G[W_{L},W_{R}]$ such that $\textup{OffDiag}(X_{L}MX_{R}^{T})$ is the adjacency matrix of $G_{\Delta}$.
Indeed, we can define $z_{i}^{L}\in\mathbb{F}_{2}^{U_{L}}$ to be the indicator vector of the edges in $G_{i}[U_{L},w^{\prime}_{i}]$ and define $z_{i}^{R}\in\mathbb{F}_{2}^{U_{R}}$ to be the indicator vector of the edges in $G_{i}[U_{R},w_{i}]$. The $i$-th pivot changes adjacency matrix of $G_{i-1}[U_{L},U_{R}]$ by exactly $z_{i}^{L}(z_{i}^{R})^{T}$. We can find upper triangular matrices $Q_{L}$ and $Q_{R}$ such that $X_{L}Q_{L}$ is the matrix with columns $(z_{i}^{L})_{i=1}^{r^{\prime}}$ and $X_{R}Q_{R}$ is the matrix with columns $(z_{i}^{R})_{i=1}^{r^{\prime}}$. Taking $M=Q_{L}Q_{R}^{T}$ satisfies the desired conditions. The rest of the argument should follow through, and we can upper bound the total variation distance between $G^{\prime}[U_{L},U_{R}]$ and $G^{\prime}(s,s,1/2)$ by $2^{-\Omega(q^{2}r^{\prime})}$ where $q=\min(p,1-p)$.
Thus, in order to prove an analogous result for pivot-minor universality, it is sufficient to answer the following question.
Question 4.1.
Let $G\sim G(r,r,p)$ and $\min(p,1-p)=\Omega(\log n/\sqrt{n})$. Can we find disjoint pairs of vertices $w_{1}w^{\prime}_{1},\dots,w_{r^{\prime}}w^{\prime}_{r^{\prime}}$ such that $w_{i+1}w^{\prime}_{i+1}$ is an edge in $G\times w_{1}w^{\prime}_{1}\times\dots\times w_{i}w^{\prime}_{i}$ for all $i=0,1,\dots,r^{\prime}-1$ and $r^{\prime}=\Theta(r)$ with high probability?
We believe that the answer to this question is affirmative since $p$ is much larger than the threshold of having a perfect matching.